OpenStudy (anonymous):

Help! Which equation could be used to solve this problem? The sum of two consecutive integers is 107. Find the integers. A. 2n + 2 = 107 B. 2n + 1 = 107 C. n + 2 = 107 D. n + 1 = 107 Which equation could be used to solve this problem? The sum of two consecutive odd integers is 112. Find the integers. A. 2n + 3 = 112 B. 2n + 2 = 112 C. 2n + 1 = 112 D. n + 2 = 112 Solve the given number problem. The sum of three consecutive even integers is –72. What is the least of these numbers? A. –26 B. –22 C. –30 D. 24

OpenStudy (anonymous):

1. Let the first integer be n then the second integer will be n+1 n+(n+1) = 107

OpenStudy (anonymous):

2. Let the first integer be n then the second odd integer will be n+2 n+(n+2) = 112

OpenStudy (anonymous):

1. Let the first integer be n the difference between three even integer is of 2 so second even integer is n+2 and third even integer is n+4 n+(n+2) + (n+4) = -72